Hi, My name is Emma, I really dont get how to do the perimeter with the equation on one side then a number with a variable next to it... like for example this square has 36-5x on one side of the square then it has 4x on the other side...please help me, I have no clue what i'm doing!
Thanks(:

Found 2 solutions by richard1234, stanbon:
Answer by richard1234(7193)

The sides of a square must all be equal, so we know that 36 - 5x = 4x. From this, we have x = 4. We can plug x = 4 to either expression (4x or 36-5x) and find that the side length of the square is 16. Since the perimeter of a square is four times one of the side lengths, the perimeter is 4*16, or 64.

Answer by stanbon(75887)

how to do the perimeter with the equation on one side then a number with a variable next to it... like for example this square has 36-5x on one side of the square then it has 4x on the other side...
---------------
The sides of a square are all equal to one another.
---
Solve the equation 36-5x = 4x
Add 5x to both sides to get:
36 = 9x
Divide both sides by 9 to solve for "x":
x = 36/9
x = 4
----
Side 36-5x = 36-5*4 = 36-20 = 16
Side 4x = 4*4 = 16
----
The dimensions of the square are 4 by 4.
=========================
Cheers,
Stan H.